How many green elements are required to make this a legitimate probability distribution if there are a total of 50 elements in this sample?
x
red
blue
orange
brown
green

P(x)
0.14(red)
0.16(blue)
0.28(orange)
0.24(brown)


Answer by ewatrrr(24785)


Hi,
P(red) + P(blue) + P(orange) + P (brown) = .14 + .16 + 28 + .24 = .82
P(green) = 1 - .82 = .18
50 elements in this sample: Let n represent the number of green elements
n/50 = .18
n = 9, the number of green elements